Hi. I am having a difficult time trying to figure out of to factor the trinomial. If you could help, that would be greatly appreciated.
2a^2 + 30a + 100. Thanks so much.

Found 2 solutions by nerdybill, Theo:
Answer by nerdybill(7384)

2a^2 + 30a + 100
Start by factoring out what is common among all three terms:
2(a^2 + 15a + 50)
now, factor what is in the parenthesis:
2(a+5)(a+10)

Answer by Theo(13342)

2a^2 + 30a + 100

set it equal to 0

you get:

2a^2 + 30a + 100 = 0

divide both sides of the equation by 2 to get:

a^2 + 15a + 50 = 0

50 can be factored in several ways.

1 * 50
2 * 25
5 * 10

looking at each of these and adding the 2 factors together you get:

1 + 50 = 51
2 + 26 = 26
5 + 10 = 15

factors of 5 and 10 look interesting because 5*10 = 50 and 5+10 = 15.

your factors could very well be:

(a + 5) * (a + 10) = 0

multiply these 2 factors together to get a^2 + 10a + 5a + 50 = 0

combine like terms to get a^2 + 15a + 50 = 0

these are your factors.

to get back to the original equation, multiply both sides of this equation by 2 to get 2a^2 + 30a + 100.

the basic procedure in factoring a quadratic equation is to get the equation into standard form and then to reduce it to the lowest possible common factors.

your expression was 2a^2 + 30a + 100.

place it into standard form by setting it equal to 0 to get:

2a^2 + 30a + 100 = 0

factor out to 2 to reduce it as much as possible to get:

a^2 + 15a + 50 = 0

etc. as we did above.

your factors came out to be (x+5) * (x+10) = 0

this equation is true if (x+5) = 0 or if (x+10) = 0.

(x+5) = 0 if x = -5
(x+10) = 0 if x = -10

your answer is:

x = -5 or x = -10